# Multi-Focused Video Group Activities Hashing

### v2

@@ 4.6 Optimization
Figure 5: M-STVH are composed of four modules: 1) Visual Module: extracting features from the input video; 2) Positional Module: modeling spatiotemporal relationships on the input position information; 3) Multi-Focused Spatiotemporal Interleaving Module: interleaving visual features and positional features at multiple layers; and 4) Hashing and Classification Learning Module: fusing the visual features with the position features and outputs the corresponding hash codes and classifications.
5 M-STVH
@@
5.1 Multi-Focused Spatiotemporal Interleaved Module
@@
Figure 6: Multi-fusion attention module, fusing visual features as well as positional features at object and group.
$\begin{array}[]{c}\boldsymbol{f}^{\prime}_{(\tau)}=\text{LayerNorm}(\boldsymbol{f}_{(\tau)}),\\ \\ \boldsymbol{f}^{\prime\prime}_{(\tau)}=\text{MFAT}(\boldsymbol{f}^{\prime}_{(\tau)},\boldsymbol{G}_{T},\boldsymbol{G}_{S}),\\ \\ \boldsymbol{f}_{(\tau+1)}=\text{FFN}(\boldsymbol{f}^{\prime\prime}_{(\tau)}),\end{array}$ (10)
@@
5.2 Hashing and Classification Learning Module
@@
$\begin{array}[]{c}\boldsymbol{h}_{(\tau)}=\boldsymbol{F}_{(\tau)}(\text{AVG}(\boldsymbol{f}_{(\tau)})),\\ \\ \boldsymbol{b}_{(\tau)}=\text{sign}(\boldsymbol{h}_{(\tau)}),\end{array}$ (12)
@@ 5.3 Loss Function
$L_{recon}=\frac{1}{d}(\boldsymbol{f}_{roi}-\boldsymbol{\widetilde{f}}_{roi})^{2},$ (13)
@@
$L_{acty}=\sum_{i=1}^{r}-w_{(i)}\boldsymbol{y}^{\text{acty }}\log\left(\widetilde{\boldsymbol{y}}_{(i)}^{\text{acty }}\right),$ (14)
@@ 5.4 Filter Matrix
$\boldsymbol{\widetilde{b}}_{(\tau)}=sign\left(\frac{\boldsymbol{b}^{\prime}_{(\tau)}-\mu}{\sigma}\right),\tau\neq\mathit{\Upsilon}$ (17)
@@
$=\frac{1}{\mathit{\Upsilon}}+\frac{1}{M\cdot(\mathit{\Upsilon}/K)},$ (18)
@@
6 Experiment
6.1 Experimental Setup
@@ 6.1.1 Datasets
Volleyball Dataset (VD) ([16]): The dataset comprises 4830 video clips extracted from 55 volleyball matches, with 3493 clips designated for training and 1337 for testing. Each clip contains 41 frames. Annotated within these clips are the coordinates of object bounding boxes, along with nine specific action labels (such as waiting, setting, digging, falling, spiking, blocking, jumping, moving, and standing) and eight group activity labels (including r-winpoint, r-passing, r-spiking, r-setting, l-setting, l-spiking, l-passing, and l-winpoint). We adhere to the actor coordinates and the training/testing partition outlined in references [47, 45, 42] to ensure comparability with group activity recognition methods.
Collective Activity Dataset (CAD) ([4]): The dataset comprises 44 video clips recorded using a low-resolution handheld camera, providing dynamic views. It encompasses five distinct collective activity labels (crossing, waiting, queueing, walking, talking), six individual action labels (NA, crossing, waiting, queueing, walking, talking), and eight individual posture labels (which are not utilized in our study). Group activity classes are determined based on the predominant actions observed within the video clip. We adopt a training/testing split of 2/3 for training and the remainder for testing to maintain consistency with prior experiments, as outlined in [47]. Additionally, following [45, 35, 44], we consolidate the classes crossing and walking into a single class moving.
Collective Activity Extended Dataset (CAED) ([5]): The CAED dataset comprises 75 video clips, wherein two additional group activity categories, dancing and jogging, have been introduced compared to the CAD dataset. Moreover, the ambiguous activity class of walking has been eliminated due to its nature as more of an individual action rather than a group activity. We adopt the training/testing partition outlined in references [47] and [43] to ensure equitable comparisons.
Table 2: The Predicted Results of Actions and Activities in VD, CAD, and CAED.’ - ’ Indicates that No Results are Provided(The best results are in bold)
@@
Contextual Model[20] RGB None - - 83.4 -
Iterativae Belief ProPagation[3] RGB None - - 79.0 -
HDTM [16] RGB AlexNet - 81.9 81.5 -
SIM[9] RGB AlexNet - - 81.2 90.2
RMIC [37] RGB+Flow Inception-v3 - 66.9 86.1 -
SBGAR [24] RGB+Flow AlexNet - - 89.4 -
SPTS [35] RGB+Flow VGG16 - 91.2 95.8 98.1
PMIC[37] RGB AlexNet - 87.7 92.2 -
ARG [40] RGB Inception-v3 83.0 92.5 91.0 -
HiGCIN [44] RGB Resnet-18 - 91.5 93.4 -
DIN [45] RGB VGG-16 - 93.6 95.9 -
P2CTDM [43] RGB Inception-v3 - 92.7 96.1 95.6
stagNet [28] RGB VGG-16 82.3 89.3 89.1 89.7
CRM [1] RGB+FLOW I3D - 93.0 85.8 -
SAVRF [27] RGB+Flow I3D 83.1 95.0 95.2 -
Dual-AI [13] RGB Inception-v3 84.4 94.4 96.5 -
Actor-Transformer [11] RGB+Flow I3D 83.7 93.0 85.8 -
GroupFormer [22] RGB Inception-v3 83.7 94.1 96.5 -
MLST-Former [47] RGB Inception-v3 84.5 94.5 96.8 95.9
RWGCN [17] RGB Resnet-50 - - 95.5 94.8
AFGNet [42] RGB Invception-v3 86.1 96.7 96.5 96.9
MOGAR [46] Keypoint HRNet - 94.5 94.4 -
@@
6.1.2 Experimental Setting
We employ the Inception-v3 model ([34]) pre-trained on ImageNet ([8]) as the CNN backbone network to ensure consistency with methods such as [35], enabling fair comparisons. For the training and testing on the VD dataset, we set $T=10$ and the video frame resolution to $H\times W=720\times 1280$. Conversely, for the CAD and CAED datasets, we also set $T=10$ but adjust the frame resolution to $H\times W=480\times 720$. Additionally, we specify the number of objects in group activities as $N=12$ for the VD dataset and $N=13$ for the CAD and CAED datasets. For group behavior retrieval, we use the training set as the base set and the testing set as the query set. For group visual analysis, we mix all videos and use video IDs as labels to evaluate the effectiveness of multi-focused video hashing retrieval. Object features are extracted using RoIAlign, which facilitates cropping and resizing to a fixed size of $5\times 5$. Optimization across all datasets is performed using Adam, with the learning rate configured as follows: for the VD dataset, the initial learning rate is set to $1\times 10^{-5}$, decaying to $5\times 10^{-6}$ at the 11th epoch and further to $1\times 10^{-6}$ at the 21st epoch. For the CAD and CAED datasets, the learning rate remains constant at $1\times 10^{-5}$ throughout the training process without decay. All datasets are trained for 60 epochs in STVH, with $\lambda_{1}$=0.1 and $\lambda_{1}$=0.5 applied uniformly. All datasets are trained for 60 epochs in M-STVH, with $\mu_{1}$=0.1, $\mu_{1}$=0.5, and $\mu_{3}$=0.01 applied uniformly. All experiments are conducted using PyTorch 1.10 with CUDA 11.3, a batch size of 2, and executed on a machine equipped with 2 GTX 4090 GPUs.
6.1.3 Evaluation Metrics
@@
6.2 Group Activity Classification
@@
As shown in Table 2, our method demonstrates strong competitiveness across multiple datasets. On the VD dataset, our approach outperforms other methods using VGG16 backbone, such as StagNet ([28]) and DIN ([45]), by nearly 3% in group activity classification. Compared with methods using Inception-v3 as the backbone network, such as MLST-Former ([47]) and GroupFormer ([22]), our method has also improved by nearly 2% in the classification of group activities, lagging only slightly behind AFGNet ([42]) by 1%. Nevertheless, our method performs classification while generating hash codes. In particular, M-STVH needs to generate multi-focused hash encoding. However, regardless of the backbone network used for feature extraction, our method achieves the best results in individual action classification. Meanwhile, on both CAD and CAED datasets, our method achieves at least 3% improvement in group activity classification accuracy compared to other state-of-the-art approaches, including AFGNet ([42]), RWGCN ([17]), and MLST-Former ([47]). Furthermore, after generating multi-focused hash encoding using M-STVH, there is no decrease in classification accuracy compared to STVH. This highlights the effectiveness of the MSF module in performing spatiotemporal interleaving at a high semantic level.
@@
Figure 7: Confusion matrix of group activity prediction on the VD, where vertical axis represent predicted labels and horizontal axis indicate ground truth
@@
Figure 8: Confusion matrix of group activity prediction on the CAD, where vertical axis represent predicted labels and horizontal axis indicate ground truth
@@
6.3 Group Activity Retrieval
@@
Table 3: Retrieved mAP on the Volleyball Dataset (best results in bold)
@@
Table 4: Retrieved mAP on the Collective Activity Dataset (best results in bold)
@@
6.4 Multi-Focused Group Activity Retrieval
@@
Figure 9: The top-1 video in multi-focus hash retrieval, where the image displayed in a green box indicates that the query sample and the retrieved sample have the same visual characteristics. In contrast, the red box indicates that they are different. Green font indicates that the query sample and the retrieved sample have the same activity category, while red indicates that they are different.
@@
Figure 10: Hierarchical retrieval performance (mAP) on VD dataset with 128-bit hash codes, evaluating both group appearance and group activity retrieval across different MSF output layers (horizontal axis)
Figure 11: In group activity inference, the attention matrices on different layers of the MSF. (a) results on VD; (b) results on CAD.
@@
Table 5: Classification accuracy and mAP@10 at 128 bits after replacing the MSF module
@@ 6.5 Ablation Experiments
6.5.1 Model Architecture Ablation Experiments
@@
Table 6: Classification Accuracy and mAP@10 at 128 bits with Different Numbers of MFS Module Stacked Layers
@@
6.5.2 Loss Function Ablation Experiments
Table 7: Classification accuracy and mAP@10 at 128 bits with different loss function constraints
@@ 6.5.2 Loss Function Ablation Experiments
Figure 12: The average Hamming distance between different class hash codes in a VD dataset at 128 bits.
As shown in Fig. 12, the comparison of average Hamming distances between hash codes reveals the effectiveness of our proposed contrastive loss. The results demonstrate that our method maintains sufficiently large Hamming distances between fundamentally different action categories (e.g.,“l-winpoint" and “r-set"), while appropriately reducing the distance between semantically similar actions (e.g., “r-set", “r-spike", and “r-pass"). The pattern indicates that the contrastive loss based on object distribution similarity successfully preserves the discriminability of dissimilar categories while bringing closer those with similar group activity patterns. The Hamming distance distribution across categories further confirms that our approach organizes the hash space in a semantically meaningful way, where the relative distances between different types of group activities correspond well to their actual behavioral relationships in volleyball games. These observations validate that the learned hash codes effectively capture both the distinctions and similarities between different group activities.
7 Conclusion
This paper presents a novel solution for group activity retrieval through two key contributions: the spatiotemporal video Hashing (STVH) model and its enhanced version, multi-focused spatiotemporal video hashing (M-STVH). The STVH framework pioneers a dual spatiotemporal interleaving approach that effectively models group activities by jointly analyzing object dynamics and group interactions, capturing both visual feature evolution and positional relationships. Building upon this foundation, M-STVH introduces a hierarchical multi-step fusion mechanism, which enables progressive integration of visual and positional features across multiple representation layers. The innovative architecture naturally transitions from shallow-layer visual feature extraction to deep-layer activity semantics understanding, while the incorporated binary filtering matrix significantly optimizes storage efficiency. The proposed method not only advances multi-focused semantic modeling in video analysis but also demonstrates substantial practical potential for applications ranging from sports analytics to intelligent surveillance systems. In future we will explore cross-camera correlation analysis to further extend the framework’s capability for large-scale scene understanding.
8 Acknowledgements
This work was supported in part by China NSF Grant No.62271274, Ningbo S&K Project Grant No. 2024Z004, No. 2023Z059, and the programs sponsored by K. C. Wong Magna Fund in Ningbo University.
9 Data Availability Statement
The datasets employed in this work are publicly accessible: the Volleyball Dataset (VD) https://doi.org/10.1109/cvpr.2016.217, the Collective Activity Dataset (CAD) https://doi.org/10.1109/iccvw.2009.5457461, and the Collective Activity Extended Dataset (CAED) https://doi.org/10.1109/cvpr.2011.5995707. All data partitions strictly follow the official evaluation protocols. The source code and implementation details for this project are publicly available at https://github.com/qzm1997/qzm1997_GAR.

### v3

@@ 4.6. Optimization
Figure 5. M-STVH are composed of four modules: 1) Visual Module: extracting features from the input video; 2) Positional Module: modeling spatiotemporal relationships on the input position information; 3) Multi-Focused Spatiotemporal Interleaving Module: interleaving visual features and positional features at multiple layers; and 4) Hashing and Classification Learning Module: fusing the visual features with the position features and outputs the corresponding hash codes and classifications.
5. M-STVH
@@
5.1. Multi-Focused Spatiotemporal Interleaved Module
@@
Figure 6. Multi-fusion attention module, fusing visual features as well as positional features at object and group.
(10) $\begin{array}[]{c}\boldsymbol{f}^{\prime}_{(\tau)}=\text{LayerNorm}(\boldsymbol{f}_{(\tau)}),\\ \\ \boldsymbol{f}^{\prime\prime}_{(\tau)}=\text{MFAT}(\boldsymbol{f}^{\prime}_{(\tau)},\boldsymbol{G}_{T},\boldsymbol{G}_{S}),\\ \\ \boldsymbol{f}_{(\tau+1)}=\text{FFN}(\boldsymbol{f}^{\prime\prime}_{(\tau)}),\end{array}$
@@
5.2. Hashing and Classification Learning Module
@@
(12) $\begin{array}[]{c}\boldsymbol{h}_{(\tau)}=\boldsymbol{F}_{(\tau)}(\text{AVG}(\boldsymbol{f}_{(\tau)})),\\ \\ \boldsymbol{b}_{(\tau)}=\text{sign}(\boldsymbol{h}_{(\tau)}),\end{array}$
@@ 5.3. Loss Function
(13) $L_{recon}=\frac{1}{d}(\boldsymbol{f}_{roi}-\boldsymbol{\widetilde{f}}_{roi})^{2},$
@@
(14) $L_{acty}=\sum_{i=1}^{r}-w_{(i)}\boldsymbol{y}^{\text{acty }}\log\left(\widetilde{\boldsymbol{y}}_{(i)}^{\text{acty }}\right),$
@@ 5.4. Filter Matrix
(17) $\boldsymbol{\widetilde{b}}_{(\tau)}=sign\left(\frac{\boldsymbol{b}^{\prime}_{(\tau)}-\mu}{\sigma}\right),\tau\neq\mathit{\Upsilon}$
@@
(18) $=\frac{1}{\mathit{\Upsilon}}+\frac{1}{M\cdot(\mathit{\Upsilon}/K)},$
@@
6. Experiment
6.1. Experimental Setup
@@ 6.1.1. Datasets
Volleyball Dataset (VD) (VD): The dataset comprises 4830 video clips extracted from 55 volleyball matches, with 3493 clips designated for training and 1337 for testing. Each clip contains 41 frames. Annotated within these clips are the coordinates of object bounding boxes, along with nine specific action labels (such as waiting, setting, digging, falling, spiking, blocking, jumping, moving, and standing) and eight group activity labels (including r-winpoint, r-passing, r-spiking, r-setting, l-setting, l-spiking, l-passing, and l-winpoint). We adhere to the actor coordinates and the training/testing partition outlined in references (MLST; DIN; AFGf) to ensure comparability with group activity recognition methods.
Collective Activity Dataset (CAD) (CAD): The dataset comprises 44 video clips recorded using a low-resolution handheld camera, providing dynamic views. It encompasses five distinct collective activity labels (crossing, waiting, queueing, walking, talking), six individual action labels (NA, crossing, waiting, queueing, walking, talking), and eight individual posture labels (which are not utilized in our study). Group activity classes are determined based on the predominant actions observed within the video clip. We adopt a training/testing split of 2/3 for training and the remainder for testing to maintain consistency with prior experiments, as outlined in (MLST). Additionally, following (DIN; SPTS; HiGCIN), we consolidate the classes crossing and walking into a single class moving.
Collective Activity Extended Dataset (CAED) (CAED): The CAED dataset comprises 75 video clips, wherein two additional group activity categories, dancing and jogging, have been introduced compared to the CAD dataset. Moreover, the ambiguous activity class of walking has been eliminated due to its nature as more of an individual action rather than a group activity. We adopt the training/testing partition outlined in references (MLST) and (P2CTDM) to ensure equitable comparisons.
Table 2. The Predicted Results of Actions and Activities in VD, CAD, and CAED.’ - ’ Indicates that No Results are Provided(The best results are in bold)
@@
Contextual Model(ContextualModel) RGB None - - 83.4 -
Iterativae Belief ProPagation(Itreative) RGB None - - 79.0 -
HDTM (VD) RGB AlexNet - 81.9 81.5 -
SIM(SIM) RGB AlexNet - - 81.2 90.2
RMIC (RMIC) RGB+Flow Inception-v3 - 66.9 86.1 -
SBGAR (SBGAR) RGB+Flow AlexNet - - 89.4 -
SPTS (SPTS) RGB+Flow VGG16 - 91.2 95.8 98.1
PMIC(RMIC) RGB AlexNet - 87.7 92.2 -
ARG (ARG) RGB Inception-v3 83.0 92.5 91.0 -
HiGCIN (HiGCIN) RGB Resnet-18 - 91.5 93.4 -
DIN (DIN) RGB VGG-16 - 93.6 95.9 -
P2CTDM (P2CTDM) RGB Inception-v3 - 92.7 96.1 95.6
stagNet (StagNet) RGB VGG-16 82.3 89.3 89.1 89.7
CRM (CRM) RGB+FLOW I3D - 93.0 85.8 -
SAVRF (SAVRF) RGB+Flow I3D 83.1 95.0 95.2 -
Dual-AI (dualAI) RGB Inception-v3 84.4 94.4 96.5 -
Actor-Transformer (Actorformer) RGB+Flow I3D 83.7 93.0 85.8 -
GroupFormer (Groupformer) RGB Inception-v3 83.7 94.1 96.5 -
MLST-Former (MLST) RGB Inception-v3 84.5 94.5 96.8 95.9
RWGCN (RWGCN) RGB Resnet-50 - - 95.5 94.8
AFGNet (AFGf) RGB Invception-v3 86.1 96.7 96.5 96.9
MOGAR (MOGAR) Keypoint HRNet - 94.5 94.4 -
@@
6.1.2. Experimental Setting
We employ the Inception-v3 model (inv3) pre-trained on ImageNet (imagenet) as the CNN backbone network to ensure consistency with methods such as (SPTS), enabling fair comparisons. For the training and testing on the VD dataset, we set $T=10$ and the video frame resolution to $H\times W=720\times 1280$. Conversely, for the CAD and CAED datasets, we also set $T=10$ but adjust the frame resolution to $H\times W=480\times 720$. Additionally, we specify the number of objects in group activities as $N=12$ for the VD dataset and $N=13$ for the CAD and CAED datasets. For group behavior retrieval, we use the training set as the base set and the testing set as the query set. For group visual analysis, we mix all videos and use video IDs as labels to evaluate the effectiveness of multi-focused video hashing retrieval. Object features are extracted using RoIAlign, which facilitates cropping and resizing to a fixed size of $5\times 5$. Optimization across all datasets is performed using Adam, with the learning rate configured as follows: for the VD dataset, the initial learning rate is set to $1\times 10^{-5}$, decaying to $5\times 10^{-6}$ at the 11th epoch and further to $1\times 10^{-6}$ at the 21st epoch. For the CAD and CAED datasets, the learning rate remains constant at $1\times 10^{-5}$ throughout the training process without decay. All datasets are trained for 60 epochs in STVH, with $\lambda_{1}$=0.1 and $\lambda_{1}$=0.5 applied uniformly. All datasets are trained for 60 epochs in M-STVH, with $\mu_{1}$=0.1, $\mu_{1}$=0.5, and $\mu_{3}$=0.01 applied uniformly. All experiments are conducted using PyTorch 1.10 with CUDA 11.3, a batch size of 2, and executed on a machine equipped with 2 GTX 4090 GPUs.
6.1.3. Evaluation Metrics
@@
6.2. Group Activity Classification
@@
As shown in Table 2, our method demonstrates strong competitiveness across multiple datasets. On the VD dataset, our approach outperforms other methods using VGG16 backbone, such as StagNet (StagNet) and DIN (DIN), by nearly 3% in group activity classification. Compared with methods using Inception-v3 as the backbone network, such as MLST-Former (MLST) and GroupFormer (Groupformer), our method has also improved by nearly 2% in the classification of group activities, lagging only slightly behind AFGNet (AFGf) by 1%. Nevertheless, our method performs classification while generating hash codes. In particular, M-STVH needs to generate multi-focused hash encoding. However, regardless of the backbone network used for feature extraction, our method achieves the best results in individual action classification. Meanwhile, on both CAD and CAED datasets, our method achieves at least 3% improvement in group activity classification accuracy compared to other state-of-the-art approaches, including AFGNet (AFGf), RWGCN (RWGCN), and MLST-Former (MLST). Furthermore, after generating multi-focused hash encoding using M-STVH, there is no decrease in classification accuracy compared to STVH. This highlights the effectiveness of the MSF module in performing spatiotemporal interleaving at a high semantic level.
@@
Figure 7. Confusion matrix of group activity prediction on the VD, where vertical axis represent predicted labels and horizontal axis indicate ground truth
@@
Figure 8. Confusion matrix of group activity prediction on the CAD, where vertical axis represent predicted labels and horizontal axis indicate ground truth
@@
6.3. Group Activity Retrieval
@@
Table 3. Retrieved mAP on the Volleyball Dataset (best results in bold)
@@
Table 4. Retrieved mAP on the Collective Activity Dataset (best results in bold)
@@
6.4. Multi-Focused Group Activity Retrieval
@@
Figure 9. The top-1 video in multi-focus hash retrieval, where the image displayed in a green box indicates that the query sample and the retrieved sample have the same visual characteristics. In contrast, the red box indicates that they are different. Green font indicates that the query sample and the retrieved sample have the same activity category, while red indicates that they are different.
@@
Figure 10. Hierarchical retrieval performance (mAP) on VD dataset with 128-bit hash codes, evaluating both group appearance and group activity retrieval across different MSF output layers (horizontal axis)
@@
Table 5. Classification accuracy and mAP@10 at 128 bits after replacing the MSF module
@@ 6.5. Ablation Experiments
6.5.1. Model Architecture Ablation Experiments
@@
Table 6. Classification Accuracy and mAP@10 at 128 bits with Different Numbers of MFS Module Stacked Layers
@@
Figure 11. In group activity inference, the attention matrices on different layers of the MSF. (a) results on VD; (b) results on CAD.
@@
6.5.2. Loss Function Ablation Experiments
Table 7. Classification accuracy and mAP@10 at 128 bits with different loss function constraints
@@ 6.5.2. Loss Function Ablation Experiments
Figure 12. The average Hamming distance between different class hash codes in a VD dataset at 128 bits.
As shown in Fig. 12, the comparison of average Hamming distances between hash codes reveals the effectiveness of our proposed contrastive loss. The results demonstrate that our method maintains sufficiently large Hamming distances between fundamentally different action categories (e.g.,“l-winpoint” and “r-set”), while appropriately reducing the distance between semantically similar actions (e.g., “r-set”, “r-spike”, and “r-pass”). The pattern indicates that the contrastive loss based on object distribution similarity successfully preserves the discriminability of dissimilar categories while bringing closer those with similar group activity patterns. The Hamming distance distribution across categories further confirms that our approach organizes the hash space in a semantically meaningful way, where the relative distances between different types of group activities correspond well to their actual behavioral relationships in volleyball games. These observations validate that the learned hash codes effectively capture both the distinctions and similarities between different group activities.
7. Conclusion
This paper presents a novel solution for group activity retrieval through two key contributions: the spatiotemporal video Hashing (STVH) model and its enhanced version, multi-focused spatiotemporal video hashing (M-STVH). The STVH framework pioneers a dual spatiotemporal interleaving approach that effectively models group activities by jointly analyzing object dynamics and group interactions, capturing both visual feature evolution and positional relationships. Building upon this foundation, M-STVH introduces a hierarchical multi-step fusion mechanism, which enables progressive integration of visual and positional features across multiple representation layers. The innovative architecture naturally transitions from shallow-layer visual feature extraction to deep-layer activity semantics understanding, while the incorporated binary filtering matrix significantly optimizes storage efficiency. The proposed method not only advances multi-focused semantic modeling in video analysis but also demonstrates substantial practical potential for applications ranging from sports analytics to intelligent surveillance systems. In future we will explore cross-camera correlation analysis to further extend the framework’s capability for large-scale
Acknowledgements.
This work was supported in part by China NSF Grant No.62271274, Ningbo S&K Project Grant No.2024Z004, No.2023Z059, Zhejiang Provincial Natural Science Foundation of China under Grant No.ZCLQN25F0207, and the programs sponsored by K. C. Wong Magna Fund in Ningbo University.